# Can We Trust AI Benchmarks? An Interdisciplinary Review of Current Issues in AI Evaluation

### v1

@@ Abstract.
Quantitative Artificial Intelligence (AI) Benchmarks have emerged as fundamental tools for evaluating the performance, capability, and safety of AI models and systems. Currently, they shape the direction of AI development and are playing an increasingly prominent role in regulatory frameworks. As their influence grows, however, so too does concerns about how and with what effects they evaluate highly sensitive topics such as capabilities, including high-impact capabilities, safety and systemic risks. This paper presents an interdisciplinary meta-review of about 100 studies that discuss shortcomings in quantitative benchmarking practices, published in the last 10 years. It brings together many fine-grained issues in the design and application of benchmarks (such as biases in dataset creation, inadequate documentation, data contamination, and failures to distinguish signal from noise) with broader sociotechnical issues (such as an over-focus on evaluating text-based AI models according to one-time testing logic that fails to account for how AI models are increasingly multimodal and interact with humans and other technical systems). Our review also highlights a series of systemic flaws in current benchmarking practices, such as misaligned incentives, construct validity issues, unknown unknowns, and problems with the gaming of benchmark results. Furthermore, it underscores how benchmark practices are fundamentally shaped by cultural, commercial and competitive dynamics that often prioritise state-of-the-art performance at the expense of broader societal concerns. By providing an overview of risks associated with existing benchmarking procedures, we problematise disproportionate trust placed in benchmarks and contribute to ongoing efforts to improve the accountability and relevance of quantitative AI benchmarks within the complexities of real-world scenarios.
@@
AI benchmarks, Benchmark critique, AI evaluation, Safety evaluation, AI Regulation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXsubmissionid: 957ccs: General and reference Evaluationccs: General and reference Metricsccs: General and reference Measurementccs: General and reference Validationccs: General and reference Reliability
1. INTRODUCTION
@@
Quantitative artificial intelligence (AI) benchmarks (i.e., combinations of test datasets and performance metrics that are taken to represent general or specific tasks and used to compare AI model capabilities and/or risks (Raji et al., 2021)) play a central role in the release and marketing of newly developed AI tools. Together with qualitative evaluation methods (such as red teaming and peer confrontations (Chollet, 2019)), quantitative benchmarks - hereafter referred to as AI benchmarks, or simply benchmarks - are generally seen as providing crucial feedback signals on the performance and capabilities of AI. Indeed, they have become so critical to AI development that businesses go to great lengths to achieve good benchmarking scores, with market players like OpenAI being estimated to have spent hundreds of thousands of dollars on compute to obtain a high score at the ARC-AGI benchmark (Pfister and Jud, 2025).
@@
Increasingly, AI benchmarks are also used in regulatory contexts, where the goal is to assess potential societal harms posed by AI models and systems. The most notable case is the EU AI Act (European Union, 2024b), which incorporates benchmarks in several key provisions. For instance, for high-risk AI systems, benchmarks are expected to play a significant role in complying with requirements on accuracy, robustness, and cybersecurity (Art. 15(2)), and their development will be facilitated through AI regulatory sandboxes (Art. 58 (2)). They will also be of fundamental importance for classifying general-purpose AI (GPAI) models with systemic risks and assessing high-impact capabilities (Art. 51(1) and Annex XIII). Furthermore, both the Board and the Scientific Panel are expected to contribute to the development of benchmarks (Art. 66(g) and Art. 68(3)). Additionally, the first and second drafts of the Code of Practice specifically mentions benchmarks as an example of best-in-class evaluations for risk assessment measures for providers of GPAI models with systemic risks (European Union, 2024a, c). In the US, benchmarks are also relevant in the recently revoked AI Executive Order (The White House, 2023) and the AI Diffusion Framework (US Department of Commerce, 2025). AI benchmarks can further be expected to play a central role in the implementation of legislations such as the EU Digital Services Act (DSA) (European Union, 2022) and the UK’s Online Safety Act (UK Parliament, 2023), which require the largest online platforms and search engines - entities that increasingly rely on AI to curate, filter, and/or rank content to millions of users - to perform regular algorithmic audits showing that their systems are safe, fair, and comply with fundamental human rights. In short, AI benchmarks - which constitute a highly heterogenous and far from standardised set of techniques - are increasingly at the heart of policy efforts to make AI more transparent and secure.
@@
At the same time however, as benchmarks are increasingly relied upon to provide AI safety assurances, researchers in a broad range of academic fields have raised serious concerns regarding their current use. This includes critical voices being raised in fields ranging from cybersecurity (McIntosh et al., 2024), linguistics (Bowman and Dahl, 2021), and computer science (Gema et al., 2024), to sociology (Engdahl, 2024), economics (Ethayarajh and Jurafsky, 2021), philosophy (LaCroix and Luccioni, 2022), ethnography (Keyes and Austin, 2022), and science and technology studies (Keyes and Austin, 2022). Such scholars have for example described current AI evaluation practices as a ”minefield” (Narayanan and Kapoor, 2023a), that raise serious ethical concerns regarding what should be measured, according to what standards, and with what downstream effects (Blili-Hamelin and Hancox-Li, 2023). They have also emphasised that benchmarks are deeply political, performative, and generative in the sense that they do not passively describe and measure how things are in the world, but actively take part in shaping it (Grill, 2024). This happens as benchmarks continuously influence how AI models are trained, fine-tuned, and applied - practices with wide-ranging political, economic, and cultural effects.
@@
With this backdrop, an interdisciplinary and up-to-date survey of quantitative AI benchmarking critique is momentously missing. The aim of this article is to address this gap by mapping and discussing known limitations in quantitative AI evaluation practices, drawing together insights from technical, cultural, and historically oriented academic fields. Our goal is to provide an overview of current risks associated with quantitative AI tests, targeting policy makers, stakeholders performing algorithmic audits, and developers of AI models and benchmarks. To achieve this, we gather and analyse around 100 publications published between 1st January 2014 and 31st December 2024 that explicitly and primarily address limitations in current benchmarking practices. Notably, more than half of these publications were published in 2023 or later, highlighting the urgency of providing an updated survey of discussions in the field.
@@
Our findings show that a rapidly growing number of researchers are voicing concerns regarding how benchmarks are used to define and measure what is safe or unsafe, moral or immoral, true or false, toxic or healthy. They further support the notion that no benchmark is neutral and that AI tests and benchmarking practices always rest ”on interwoven technical and normative decisions” (Rauh et al., 2024, p. 1201) in ways that urge policy makers and AI developers to apply them with caution. In particular, previous research highlights a need to question the relevance and trustworthiness of well-cited benchmarks, since existing studies have repeatedly shown weaknesses in benchmarks perceived as state-of-the-art (SOTA). Our findings also reveal a substantial need to scrutinise capability and safety-oriented AI benchmarks to the same extent as the AI models they are meant to evaluate. In short, AI benchmarks need to be subjected to the same demands concerning transparency, fairness, and explainability, as algorithmic systems and AI models writ large.
@@
In what follows, we first provide a background discussion on AI benchmarking practices, including clarifications on the terminology used in this paper. Next, we situate our work within a series of previous research that summarise benchmark critique, before providing a summary of nine issues with quantitative benchmarks, identified during the course of our research. While not exhaustive, we present these issues as a taxonomy of benchmark critique that highlights crucial and often interrelated points of concern regarding benchmarks, voiced in the past decade. Finally, we provide some concluding remarks on the implications these issues have for policy makers and implementers. On the whole, our work constitutes the first step in a broader research and policy-oriented project aimed at developing a framework for trustworthy AI benchmarks, and will proceed by considering mitigation strategies.
2. BACKGROUND
@@
Etymologically, the term ”benchmark” has its roots in land surveying, where a physical mark (known as a ‘bench mark’) was used as a reference point for measuring elevations. This mark typically consisted of a horizontal groove in a surface, which supported a level surface or ’bench’ for a levelling rod. Over time, the term evolved to encompass a broader meaning, referring to any standard or reference point used for comparison or evaluation (Oxford English Dictionary, 2017). Benchmarking is currently applied across many different domains such as bioinformatics (Aniba et al., 2010), environmental quality (Henning, 2000), information retrieval (Thakur et al., 2021), transistor hardware (Cheng et al., 2020), industry and business (Camp, 1989), and within the public sector (Bruno, 2014), where it often refers to procedures for comparing the performance or best practices of different companies or processes. Here, we focus on the use of benchmark tests within computing, where they are used to evaluate the performance of hardware or software systems by comparing them to a standard or reference point (Henning, 2000). More specifically, we zoom in on AI development, where benchmarks are often used to facilitate cross-model comparisons, measure performance, track model progress, and identify weaknesses (Reuel et al., 2024). Benchmarks are often perceived as comparably cost- and time-effective tools for AI providers who may run them regularly throughout the model development to obtain frequent signals of AI model performance and capabilities (Weidinger et al., 2023). They can be applied to both software and hardware solutions, where the latter for instance evaluates the performance of CPUs, GPUs, TPUs (Mattson et al., 2020b), or hardware accelerators (Mattson et al., 2020a). We address software-oriented benchmarks. Based on the definition proposed by Raji et al. (2021), we define such benchmarks as “a particular combination of a set of test datasets, including human-in-the-loop interactions, and associated performance metrics, conceptualised as representing one or more specific tasks or related capabilities, typically chosen by a community of researchers as a shared framework for comparing AI models”. A testing dataset would refer to “a separate dataset, distinct from the training data, used to objectively evaluate the performance and generalisability of an AI model”. These datasets are typically composed of samples or instances that include an input paired with the desired output (also known as reference, gold, or ground truth). A task would be “a particular specification of a problem, typically represented as a mapping between an input space and an output or action space, extensionally defined through one or more datasets, and usually associated with a specific performance evaluation metric” (based on (Raji et al., 2021) and (Schlangen, 2020)). And finally, a metric would refer to ”a specification of the mechanism to determine the degree of success or failure of the model’s outputs”. The metric represents a way to summarise model performance over a given task and dataset, usually defined as a single number or score (Raji et al., 2021). AI models that achieve the best scores on the metrics of a benchmark are generally considered SOTA in terms of performance.
@@
While humans are always involved in the design and creation of ground truth in benchmarks, they can play a more or less direct and active role in the process of applying benchmark tests. In automated or quantitative benchmarks, a set of tasks, datasets, and metrics are first defined through human decision making. The execution of a benchmark test is then carried out without direct human intervention. In qualitative benchmarks, humans directly intervene and partake in evaluations, for example as evaluators, judges, or real-time interrogators (e.g., adversarial testing or red teaming). We primarily consider quantitative benchmarks. This is not to say that there is no relevant critique concerning qualitative AI, but such methods introduce a different set of critique that is out of scope for our discussions here.
3. RELATED WORK
@@
Over the years, several surveys and meta-reviews have set out to summarize discussions on limitations in AI benchmarking and this paper represents a continuation of such efforts. In 2021, Liao et al. (2021, p. 1) identified a wide range of ”surprisingly consistent critique” directed at benchmarking practices across fields such as computer vision, natural language processing, recommender systems, reinforcement learning, graph processing, metric learning, and more. The authors present a taxonomy of observed benchmark failure modes, which for example includes implementation variations (due to benchmark algorithms, metrics, and libraries being applied in slightly different ways), errors in test set construction (as a result of label errors, label leakage, test set size, or contaminated data), overfitting from test set reuse, and comparisons to inadequate baselines. Along similar lines, Hutchinson et al. (2022) survey discussions on evaluation practices in machine learning, particularly in the fields of computer vision and natural language processing. The authors identify eight key evaluation gaps, including topics such as ”neglect of model interpretability” and ”oversimplification of knowledge” and argue for a shift toward application-centric evaluations that account for safety, fairness, and ecological validity. Similar points are also raised by Gehrmann et al. (2023)’s most recent meta-review, which categorizes issues with text-oriented benchmarks identified over the past two decades, and propose a long-term vision for improving evaluation practices, emphasizing the need for comprehensive evaluation reports that include multiple datasets, metrics, and human assessments. Their key recommendations include focusing on model limitations, enhancing dataset documentation, and adopting a more nuanced approach to evaluation to better characterize model performance and capabilities.
@@
Our review revisits these areas of critique and traces how debates concerning benchmarks have evolved in the last few years. This is not least necessary due to rapid developments currently seen in the field. Recent research has identified an unprecedented growth in the release of AI benchmarks, especially in the area of safety, starting from 2023 and onwards (Röttger et al., 2024). Our research indicates a similar increase in publications expressing benchmark critique. For instance, Röttger et al. (2024) found that roughly 46% of the AI safety benchmarks they identified as relevant in their survey of the field had been produced in 2023, with as many as 15 new datasets being released during the first two months of 2024. Likewise, almost 55% of the articles we identified as relevant during the course of our research had been released in 2023 or later. Given this rapid increase in attention given to the topic, our primary goal is to extend existing meta-reviews into the present. Whereas previous meta-reviews have tended to focus on critique raised in a limited number of domains (primarily natural language processing and machine learning) we also include interdisciplinary accounts from the humanities and social sciences in our review. Furthermore, we provide insights concerning more recently identified areas of benchmark concern, such as works revealing that AI models can be programmed to underperform on benchmarking tests (a problem known as ”sandbagging”) (Weij et al., 2024), and studies presenting evidence that AI safety benchmarks strongly correlate with upstream AI capabilities (causing reasons to worry about AI ”safetywashing”) (Ren et al., 2024), insights that both raise additional questions about the validity and trustworthiness of current benchmarking practices.
4. METHODOLOGY
Since benchmarks constitute a key element in most papers that discuss or demonstrate the capabilities of AI models, it is challenging to identify previous works that explicitly and thoroughly address benchmark critique, as opposed to applying benchmarks or proposing new methods. When gathering source materials for this review (and instead of conducting keyword searches in research databases) we therefore opted for a method that involved starting from a selection of well-cited papers that problematise how, when, and where benchmarks are used. For each article, we surveyed its reference list and studied how the paper has been cited since its publication. This allowed us to identify additional relevant literature and expand our corpus successively.

### v2

@@ Abstract
Quantitative Artificial Intelligence (AI) Benchmarks have emerged as fundamental tools for evaluating the performance, capability, and safety of AI models and systems. Currently, they shape the direction of AI development and are playing an increasingly prominent role in regulatory frameworks. As their influence grows, however, so too does concerns about how and with what effects they evaluate highly sensitive topics such as capabilities, including high-impact capabilities, safety and systemic risks. This paper presents an interdisciplinary meta-review of about 100 studies that discuss shortcomings in quantitative benchmarking practices, published in the last 10 years. It brings together many fine-grained issues in the design and application of benchmarks (such as biases in dataset creation, inadequate documentation, data contamination, and failures to distinguish signal from noise) with broader sociotechnical issues (such as an over-focus on evaluating text-based AI models according to one-time testing logic that fails to account for how AI models are increasingly multimodal and interact with humans and other technical systems). Our review also highlights a series of systemic flaws in current benchmarking practices, such as misaligned incentives, construct validity issues, unknown unknowns, and problems with the gaming of benchmark results. Furthermore, it underscores how benchmark practices are fundamentally shaped by cultural, commercial and competitive dynamics that often prioritise state-of-the-art performance at the expense of broader societal concerns.
By providing an overview of risks associated with existing benchmarking procedures, we problematise disproportionate trust placed in benchmarks and contribute to ongoing efforts to improve the accountability and relevance of quantitative AI benchmarks within the complexities of real-world scenarios.
@@
Keywords AI Benchmarks  $\cdot$ Benchmark Critique  $\cdot$ AI Evaluation  $\cdot$ Safety Evaluation  $\cdot$ AI Regulation
1 Introduction
@@
Quantitative artificial intelligence (AI) benchmarks (i.e., combinations of test datasets and performance metrics that are taken to represent general or specific tasks and used to compare AI model capabilities and/or risks [Raji et al., 2021]) play a central role in the release and marketing of newly developed AI models. Together with qualitative evaluation methods (such as red teaming and peer confrontations [Chollet, 2019]), quantitative benchmarks - hereafter referred to as AI benchmarks, or simply benchmarks - are generally seen as providing crucial feedback signals on the performance and capabilities of AI. Indeed, they have become so critical to AI development that businesses go to great lengths to achieve good benchmarking scores, with market players like OpenAI being estimated to have spent hundreds of thousands of dollars on compute to obtain a high score at the ARC-AGI benchmark [Pfister and Jud, 2025].
@@
Increasingly, AI benchmarks are also used in regulatory contexts, where the goal is to assess potential societal harms posed by AI models and systems. The most notable case is the EU AI Act [European Union, 2024c], which incorporates benchmarks in several key provisions such as high-risk assessments of AI systems, where they are expected to inform requirements on accuracy, robustness, and cybersecurity (Art. 15(2)), facilitated through AI regulatory sandboxes (Art. 58 (2)). They will also be of fundamental importance for classifying general-purpose AI (GPAI) models with systemic risks and assessing high-impact capabilities (Art. 51(1) and Annex XIII). Additionally, the first and second drafts of the Code of Practice for the EU AI Act mentions benchmarks as a risk assessment method for providers of GPAI models with systemic risks [European Union, 2024a, b]. In the US, benchmarks are also relevant in the recently revoked AI Executive Order [The White House, 2023] and the AI Diffusion Framework [US Department of Commerce, 2025]. AI benchmarks can further be expected to play a central role in the implementation of legislations such as the EU Digital Services Act (DSA) [European Union, 2022] and the UK’s Online Safety Act [UK Parliament, 2023], which require the largest online platforms and search engines - entities that increasingly rely on AI to curate, filter, and/or rank content to millions of users - to perform regular algorithmic audits. In short, AI benchmarks - which constitute a highly heterogenous and far from standardised set of techniques - are increasingly at the heart of policy efforts to make AI more transparent and secure.
@@
At the same time as benchmarks are increasingly relied upon to provide AI safety assurances, however, journalists [Keegan, 2024; Roose, 2024; Heaven, 2023] and researchers in a broad range of academic fields have raised serious concerns regarding their current use. This includes critical voices being raised in fields ranging from cybersecurity [McIntosh et al., 2024], linguistics [Bowman and Dahl, 2021], and computer science [Gema et al., 2024], to sociology [Engdahl, 2024], economics [Ethayarajh and Jurafsky, 2021], philosophy [LaCroix and Luccioni, 2022], ethnography [Keyes and Austin, 2022], and science and technology studies [Keyes and Austin, 2022]. Such scholars have for example described current AI evaluation practices as a "minefield" [Narayanan and Kapoor, 2023b], that raise serious ethical concerns regarding what should be measured, according to what standards, and with what downstream effects [Blili-Hamelin and Hancox-Li, 2023; Mitchell, 2023; Burden, 2024]. They have also emphasised that benchmarks are deeply political, performative, and generative in the sense that they do not passively describe and measure how things are in the world, but actively take part in shaping it [Grill, 2024]. This happens as benchmarks continuously influence how AI models are trained, fine-tuned, and applied - practices with wide-ranging political, economic, and cultural effects.
@@
With this backdrop, an interdisciplinary and up-to-date survey of quantitative AI benchmarking critique is momentously missing. The aim of this article is to address this gap by mapping known limitations in quantitative AI evaluation practices, drawing together insights from multiple academic fields. Our goal is to provide an overview of current risks associated with quantitative AI tests, targeting policy makers, stakeholders performing algorithmic audits, and developers of AI models and benchmarks. To achieve this, we gather and analyse around 110 publications published between 1st January 2014 and 31st December 2024 that explicitly and primarily address limitations in current benchmarking practices. Notably, more than half of these publications were published in 2023 or later, highlighting the urgency of providing an updated survey of discussions in the field.
@@
Our findings show that a rapidly growing number of researchers are voicing concerns regarding how benchmarks are used to define and measure what is safe or unsafe, moral or immoral, true or false, toxic or healthy. They further support the notion that no benchmark is neutral and that AI tests and benchmarking practices always rest "on interwoven technical and normative decisions" [Rauh et al., 2024, p. 1201] in ways that urge policy makers and AI developers to apply them with caution. In particular, previous research highlights a need to question the relevance and trustworthiness of well-cited benchmarks, since existing studies have repeatedly shown weaknesses in benchmarks perceived as state-of-the-art (SOTA). Our findings also reveal a substantial need to scrutinise capability and safety-oriented AI benchmarks to the same extent as the AI models they are meant to evaluate. In short, AI benchmarks need to be subjected to the same demands concerning transparency, fairness, and explainability, as algorithmic systems and AI models writ large.
@@
In what follows, we first provide a background discussion on AI benchmarking practices, including clarifications on the terminology used in this paper. Next, we situate our work within a series of previous research that summarise benchmark critique, before providing a summary of nine issues with quantitative benchmarks, identified during the course of our research. While not exhaustive, we present these issues as a taxonomy of benchmark critique that highlights crucial and often interrelated points of concern regarding benchmarks, voiced in the past decade. Finally, we provide some concluding remarks on the implications these issues have for policy makers. On the whole, our work constitutes the first step in a broader research and policy-oriented project aimed at developing a framework for trustworthy AI benchmarks, and will proceed by considering mitigation strategies.
2 Background
@@
Etymologically, the term “benchmark” has its roots in land surveying, where a physical mark (known as a ‘bench mark’) was used as a reference point for measuring elevations. This mark typically consisted of a horizontal groove in a surface, which supported a level surface or ’bench’ for a levelling rod. Over time, the term evolved to encompass a broader meaning, referring to any standard or reference point used for comparison or evaluation [Oxford English Dictionary, 2017]. Benchmarking is currently applied across many different domains such as bioinformatics [Aniba et al., 2010], environmental quality [Henning, 2000], security [Noroozian, 2020], information retrieval [Thakur et al., 2021], transistor hardware [Cheng et al., 2020], industry and business [Camp, 1989], and within the public sector [Bruno, 2014], where it often refers to procedures for comparing the performance or best practices of different companies or processes. Here, we focus on the use of benchmark tests within computing, where they are used to evaluate the performance of hardware or software systems by comparing them to a standard or reference point [Henning, 2000]. More specifically, we zoom in on AI development, where benchmarks often facilitate cross-model comparisons, track model progress, and identify weaknesses [Reuel et al., 2024].
Benchmarks are often perceived as comparably cost- and time-effective tools for AI providers who may run them regularly throughout the model development to obtain signals of model capabilities [Weidinger et al., 2023]. They can be applied to both software and hardware solutions, where the latter, for instance, evaluates the performance of CPUs, GPUs, or TPUs [Mattson et al., 2020b], and/or hardware accelerators [Mattson et al., 2020a]. We address software-oriented benchmarks. Based on the definition proposed by Raji et al. [2021], we define such benchmarks as “a particular combination of a set of test datasets, including human-in-the-loop interactions, and associated performance metrics, conceptualised as representing one or more specific tasks or related capabilities, typically chosen by a community of researchers as a shared framework for comparing AI models”. A testing dataset would refer to “a separate dataset, distinct from the training data, used to objectively evaluate the performance and generalisability of an AI model”. These datasets are typically composed of samples or instances that include an input paired with the desired output (also known as reference, gold, or ground truth). A task would be “a particular specification of a problem, typically represented as a mapping between an input space and an output or action space, extensionally defined through one or more datasets, and usually associated with a specific performance evaluation metric” (based on [Raji et al., 2021] and [Schlangen, 2020]). And finally, a metric would refer to "a specification of the mechanism to determine the degree of success or failure of the model’s outputs". The metric represents a way to summarise model performance over a given task and dataset, usually defined as a single number or score [Raji et al., 2021]. AI models that achieve the best scores on the metrics of a benchmark are generally considered SOTA in terms of performance.
@@
While humans are always involved in the design and creation of ground truth in benchmarks, they can play a more or less direct and active role in applying benchmark tests. In automated or quantitative benchmarks, a set of tasks, datasets, and metrics are first defined through human decision making. The execution of a benchmark test is then carried out without direct human intervention. In qualitative benchmarks, humans intervene and partake in evaluations, for example, as evaluators, judges, or real-time interrogators (e.g., adversarial testing or red teaming). We primarily consider quantitative benchmarks. This is not to say that there is no relevant critique concerning qualitative AI evaluations, but such methods introduce a different set of problems that is out of scope for our discussions here.
3 Related Work
@@
Over the years, several surveys and meta-reviews have set out to summarize discussions on limitations in AI benchmarking and this paper represents a continuation of such efforts. In 2021, Liao et al. [2021, p. 1] identified a wide range of "surprisingly consistent critique" directed at benchmarking practices across fields such as computer vision, natural language processing, recommender systems, reinforcement learning, graph processing, metric learning, and more. The authors present a taxonomy of observed benchmark failure modes, including implementation variations, errors in test set construction, overfitting from test set reuse, and comparisons to inadequate baselines. Along similar lines, Hutchinson et al. [2022] survey discussions on evaluation practices in machine learning, particularly in the fields of computer vision and natural language processing. The authors identify eight key evaluation gaps, including topics such as "neglect of model interpretability" and "oversimplification of knowledge" and argue for a shift toward application-centric evaluations that account for safety, fairness, and ecological validity. Similar points are also raised by Gehrmann et al. [2023]’s most recent meta-review, which categorizes issues with text-oriented benchmarks over the past two decades, and propose a long-term vision for improving evaluation practices, emphasizing the need for comprehensive evaluations that include multiple datasets, metrics, and human assessments. Their key recommendations include focusing on model limitations, enhancing dataset documentation, and adopting a more nuanced approach to evaluation to better characterize model capabilities.
@@
Our review revisits these areas of critique and traces how debates concerning benchmarks have evolved in the last few years due to rapid developments in the field. Recent research has identified an unprecedented growth in the release of AI benchmarks, especially in the area of safety, starting from 2023 and onwards [Röttger et al., 2024]. Our research indicates a similar increase in publications expressing benchmark critique. For instance, Röttger et al. [2024] found that roughly 46% of the AI safety benchmarks they identified as relevant in their survey of the field had been produced in 2023, with as many as 15 new datasets being released during the first two months of 2024. Likewise, almost 55% of the articles we identified as relevant during the course of our research had been released in 2023 or later. Given this rapid increase in attention given to the topic, our primary goal is to extend existing meta-reviews into the present. Whereas previous meta-reviews have tended to focus on critique raised in a limited number of domains (primarily natural language processing and machine learning) we also include interdisciplinary accounts from the humanities and social sciences in our review. Furthermore, we provide insights concerning more recently identified areas of benchmark concern, such as works revealing that AI models can be programmed to underperform on benchmarking tests (a problem known as “sandbagging”) [Weij et al., 2024], and studies presenting evidence that AI safety benchmarks strongly correlate with upstream AI capabilities (causing reasons to worry about AI “safetywashing”) [Ren et al., 2024], insights that both raise additional questions about the validity and trustworthiness of current benchmarking practices.
4 Methodology
When gathering source materials for this review (and instead of conducting systematic keyword searches in research databases) we opted for a snowball sample [Jalali and Wohlin, 2012; Badampudi et al., 2015], which helped us identify papers that primarily address benchmark critique (our main inclusion criteria) while avoiding noise in the form of articles that present new benchmarks or simply apply them (our main exclusion criteria). A snowball method also helped us cut across interdisciplinary inconsistencies in terminology and allowed us to discover pre-prints - a publication form where much debate concerning benchmarks has taken place, but that would have been missed in a traditional review based on scientific databases. We initially started from the article "AI and the Everything in the Whole Wide World Benchmark" [Raji et al., 2021] and expanded by reviewing articles included in its reference list. We also studied how the text has been cited since its publication using Google Scholar - a process that was repeated for each article considered relevant. When important concepts (such as sandbagging) appeared in the literature, targeted searches for research in these fields were made. We also made special efforts to include discussions concerning benchmarks for images, sound, and audiovisual content which are under-represented compared to text-focused benchmarks.